Find the strength of electric field at distance of 0.5m from a point of 10uc situated in a medium of dielectric constant 5.


1
Expert's answer
2021-03-18T14:03:55-0400
"E=k\frac{q}{\epsilon r^2}\\E=9\cdot10^{9}\frac{10\cdot10^{-6}}{(5)(0.5)^2}=72000\frac{N}{C}"